# positioning

### v1

@@ 6.2. The main statement
has a natural embedding in $\mathcal{O}_{K}$, the ring of integers of a totally real number field $K$. There are two ways to show that $\mathbb{Z}(\rho)$ lies in $\mathcal{O}_{K}$, as opposed to the ring of $S$-integers for some finite set of finite places $S$ of $K$. In [8, Sec. 4] it is done by invoking the recent work of Esnault and Groechenig [17]. However, here we can give a self-contained proof by applying 1.12.
From here forward we assume that $\rho(\Gamma)$ is Zariski dense in ${\mathbf{H}}(k)$ and will then prove that ${\mathbf{H}}(k)\cong G$. In particular, from the above we obtain a natural $K$-form $\mathbf{H}\subset$ of ${\mathbf{H}}$. Let $\sigma_{i}:K\to\mathbb{R}$ be the places of $K$, ordered in such a way that $\sigma_{1}$ is the identity for the given lattice embedding. Let $\widehat{{\mathbf{H}}}\subset=\operatorname{Res}_{K/\mathbb{Q}}\mathbf{H}\subset$ and consider the $\mathbb{Z}$-local system
$\hat{\rho}:\Gamma$ $\longrightarrow\widehat{\mathbf{H}}\subset(\mathbb{R})$
@@
$\𝐷\psi:S_{\Gamma}^{\textnormal{an}}\longrightarrow\widehat{\mathbf{H}}\subset(\mathbb{Z})\backslash D.$
Now that we have a $\mathbb{Z}$VHS, as in 6.10 we use the results of Section 6 to prove the theorem. Let $S_{\Gamma}=\Gamma\backslash\mathbb{B}^{n}$ be the associated ball quotient, and $\mathbb{V}_{1}$ the $\mathbb{Z}$VHS on $S_{\Gamma}$ (associated to some faithful linear rational representation of $\mathbf{G}$). By construction $\textnormal{HL}(S_{\Gamma},\mathbb{V}_{1}^{\otimes})_{\textnormal{pos}}$ is the union of the complex totally geodesic subvarieties of $S_{\Gamma}$, and it is equal to $\textnormal{HL}(S_{\Gamma},\mathbb{V}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$ (in particular it is non-empty). It is straightforward to see that the fact that $\rho$ is complex rich implies that $\hat{\rho}$ is complex rich as well. In particular, it follows that
(7.0.1) $\textnormal{HL}(S_{\Gamma},\mathbb{V}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}=\textnormal{HL}(S_{\Gamma},\widehat{\mathbb{V}}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$
6.4 implies that $\mathbb{V}_{1},\widehat{\mathbb{V}}$ are isogenous which means that the period domain $\𝐷\widehat{\mathbf{H}}\subset(\mathbb{Z})\backslash D$ is isomorphic to $S_{\Gamma}$, and that $\psi$ is isogenous to the identity (as in 6.3). Therefore $\hat{\rho}$ extends, which is only possible if the initial representation $\rho$ extends. That is, $\psi$ is an isomorphism. This concludes the proof of the desired statement. ∎
Proof of 1.14.
The proof follows the same argument given above. The only difference is we are given a $\mathbb{Q}$VHS ${\mathbb{V}}$ on the algebraic variety $S_{\Gamma}$. By applying 1.12, ${\mathbb{V}}$ must be a $\mathbb{Z}$VHS, and the richness assumption implies that
$\textnormal{HL}(S_{\Gamma},{\mathbb{V}}^{\otimes})_{\textnormal{pos},\textnormal{typ}}=\textnormal{HL}(S_{\Gamma},{\mathbb{V}}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}\neq\emptyset,$
as in (7.0.1) above. Therefore the results follows, once more, from 6.4. ∎

### v3

@@ 6.2. The main statement
has a natural embedding in $\mathcal{O}_{K}$, the ring of integers of a totally real number field $K$. There are two ways to show that $\mathbb{Z}(\rho)$ lies in $\mathcal{O}_{K}$, as opposed to the ring of $S$-integers for some finite set of finite places $S$ of $K$. In [8, Sec. 4] it is done by invoking the recent work of Esnault and Groechenig [18]. However, here we can give a self-contained proof by applying 1.14.
From here forward we assume that $\rho(\Gamma)$ is Zariski dense in the $k$-group ${\mathbf{H}}$ and will then prove that ${\mathbf{H}}(K)\cong G$. In particular, from the above we obtain a natural $K$-form of ${\mathbf{H}}$ (that we denote by the same symbol ${\mathbf{H}}$). Let $\sigma_{i}:K\to\mathbb{R}$ be the places of $K$, ordered in such a way that $\sigma_{1}$ is the identity for the given lattice embedding. Let $\widehat{{\mathbf{H}}}\subseteq\operatorname{Res}_{K/\mathbb{Q}}\mathbf{H}$ and consider the $\mathbb{Z}$-local system
$\hat{\rho}:\Gamma$ $\longrightarrow\widehat{\mathbf{H}}(\mathbb{R})$
@@
$\psi:S_{\Gamma}^{\textnormal{an}}\longrightarrow\widehat{\mathbf{H}}(\mathbb{Z})\backslash D.$
Now that we have a $\mathbb{Z}$VHS, as in 6.10 we use the results of Section 6 to prove the theorem. Let $S_{\Gamma}=\Gamma\backslash\mathbb{B}^{n}$ be the associated ball quotient, and $\mathbb{V}_{1}$ the $\mathbb{Z}$VHS on $S_{\Gamma}$ associated with some faithful linear rational representation of $\mathbf{G}$. By construction, $\textnormal{HL}(S_{\Gamma},\mathbb{V}_{1}^{\otimes})_{\textnormal{pos}}$ is the union of the complex totally geodesic subvarieties of $S_{\Gamma}$, and it is equal to $\textnormal{HL}(S_{\Gamma},\mathbb{V}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$; in particular it is nonempty.
It is straightforward to see that the fact that $\rho$ is complex rich implies that $\hat{\rho}$ is complex rich as well. Indeed, $S_{\Gamma_{i}}\subset S_{\Gamma}$ be a totally geodesic subvariety such that $\mathbf{H}_{i}=\overline{\rho(\Gamma_{i})}$ has dimension smaller than $\dim\mathbf{H}$. Then $\dim\widehat{\mathbf{H}}_{i}:=\overline{\hat{\rho}(\Gamma_{i})}<\dim\widehat{\mathbf{H}}$. Moreover $S_{\Gamma_{i}}$ lies in the $\textnormal{HL}(S_{\Gamma},\mathbb{V}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$, as well as $\textnormal{HL}(S_{\Gamma},\widehat{\mathbb{V}}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$. Since this is true for a sequence of $i$ (the proof of) 6.4 implies that $\mathbb{V}_{1},\widehat{\mathbb{V}}$ are isogenous which means that the period domain $\widehat{\mathbf{H}}(\mathbb{Z})\backslash D$ is isomorphic to $S_{\Gamma}$ and that $\psi$ is isogenous to the identity (as in 6.3). Therefore $\hat{\rho}$ extends, which is only possible if the initial representation $\rho$ extends. That is, $\psi$ is an isomorphism. This concludes the proof of the desired statement. ∎
Proof of 1.17.
The proof follows the same argument given above. The only difference is we are given a $\mathbb{Q}$VHS ${\mathbb{V}}$ on the algebraic variety $S_{\Gamma}$. By applying 1.14, ${\mathbb{V}}$ must be a $\mathbb{Z}$VHS, and the richness assumption implies that there is a generic sequence of subvarieties $S_{\Gamma_{i}}\in\textnormal{HL}(S_{\Gamma},{\mathbb{V}}^{\otimes})_{\textnormal{pos},\textnormal{typ}}\cap\textnormal{HL}(S_{\Gamma},{\mathbb{V}}_{1}^{\otimes})_{\textnormal{pos},\textnormal{typ}}$. Therefore the results follows once again from 6.4. ∎